OpenStudy (rock_mit182):

Suppose that the relation G is defined as follows. G = {( 5 ,0 ), (7,-7 ), (-8,-7 ) } Give the domain and range of . Write your answers using set notation

OpenStudy (rock_mit182):

DOMAIN { 5,7,-8} RANGE {0,-7,-7 } right ?

OpenStudy (anonymous):

You got it :)

OpenStudy (rock_mit182):

r (x ) = x+3 ; s(X) = 4x^2 (r * s) (x)= 4x^3 +12x^2

OpenStudy (rock_mit182):

I hope is right...

OpenStudy (rock_mit182):

(r-s) = -4x^2 +x+3 (r+s)(-1) = 4(-1)^ + -1 +3

OpenStudy (rock_mit182):

(r+s)(-1) = 6